# GravSphere2: A higher-order Jeans method for mass-modeling spherical stellar systems

### v2

@@ 4.3 Low tracer numbers and comparisons with mass estimators
Some of the most interesting objects to study at any given time are either extremely intrinsically faint or distant, meaning that they are (and perhaps always will be) scarce in tracer numbers. [^21] This should not be interpreted as prognosticating pessimism for future observations, but rather that with increasing observations it appears likely that interest will shift towards ever fainter systems that lie at the limit of the capabilities at any given time. Some current examples include the discovery of the faintest known Milky Way satellite (2024ApJ...961...92S) or the dwarf galaxies in Andromeda (2021MNRAS.505.5686C; 2025arXiv250504475P). In these cases, one is limited to, at most, a few tens of kinematic tracers. It is therefore of considerable interest to study the behavior of GravSphere2 in this low tracer limit, comparing it with other commonly-applied methods such as the mass estimators from 2009ApJ...704.1274W; 2010MNRAS.406.1220W; 2017MNRAS.469.2335C; 2018MNRAS.481.5073E. We have done this for the case of the Fornax simulated dwarf, showing our results in Fig. 7 (left). [^22] We have omitted showing the mass estimator result from 2017MNRAS.469.2335C as it produced a very similar result that overlapped with the one from 2018MNRAS.481.5073E. For this case, we computed the global velocity dispersion by fitting GravSphere2’s non-Gaussian PDFs from Sec. 2.2 (i.e. creating one bin for all the tracers), although Gaussian PDF instead did not change our results (in this case). The errors on the estimator were then propagated from the posterior distribution of the global dispersion directly, assuming no other sources of error (systematic or otherwise). In this case, all estimators show good agreement with GravSphere2, albeit with somewhat larger errors, with all recovering the true solution within the specified 95% CL regions.
@@
Therefore, while the nominal errors of mass estimators are comparable to those of GravSphere2, the simple estimators suffer from potentially significant systematic biases that are not present in GravSphere2. We conclude that even for very low numbers of tracers – of order 10 – it is still worth using GravSphere2 rather than these estimators, which has the additional advantage of giving constraints on the mass profile at all radii, rather than just one point at $\sim R_{1/2}$. Alternatively, particularly for more robust choices like the Errani estimator, biases can be substantially reduced when adopting non-Gaussian PDFs, but these can still creep in as the number of tracers is increased.
Note that in our main analysis, we have assumed a good knowledge of the photometric tracer profile, using our default value of 10,000 tracers for most tests so far. While it is often the case that the photometric light profile is far better constrained than the kinematic one, one may ask what the effects of restricting ourselves to a low number of photometric tracers might be on the mass estimates. Interestingly, repeating our analysis using the kinematic tracers as the only source of photometric information yielded essentially identical results. This indicates that the advantage of GravSphere2 over the mass estimators does not stem from additional information provided by a good knowledge of the tracer profile, and that the uncertainty in the kinematic information dominates the uncertainty in our mass estimates.
@@ 5 Discussion and conclusion
In this paper, we have introduced and tested GravSphere2, an extended Jeans modeling tool that includes full treatment of the higher-order Jeans equations of fourth-order moments. We have introduced fourth-order PMs as a new observable to break model degeneracies, using a general radially dependent fourth-order anisotropy term as part of our analysis. This extends previous works, which have only considered higher moments for line-of-sight velocities and/or made restrictive assumptions about the form of the higher-order counterpart of the anisotropy.
@@
Unlike previous versions of GravSphere, our methodology does not rely on VSPs which can become biased (see discussion at the end of Sec. 2.1). It is also completely ‘bin-free’, with a generalized velocity PDF that we use to fit discrete data star-by-star. GravSphere2 includes an improved tracer density model that has the option of fitting tracer positions individually, and an improved sampling technique. GravSphere2 does not rely on assumptions about the form of the phase-space distribution function, as it implicitly constrains it through its velocity moments via the second and fourth order Jeans equations. In this way, our method retains the flexibility and bias-free nature of usual Jeans analyses by not needing to specify any particular form for the distribution function, while gaining constraining power from higher-order velocity information, avoiding negative moments that are unphysical (at fourth order), and having all the other general advantages of having a distribution function.
@@
We tested these improvements using equilibrium mock data from the Gaia Challenge suite, comparing our results to those from 2021MNRAS.501..978R and 2021MNRAS.505.5686C, where multiple methods were tested, including GravSphere1.5, the (until now) latest version of GravSphere, Agama, a popular distribution function-based method, and DiscreteJAM, an axisymmetric Jeans-based method. We also tested our method with simulated dwarf galaxy data recently presented in 2025arXiv250418617T, where PyGravSphere (that we refer to in this paper as GravSphere1) was also tested.
@@
We also tested GravSphere2 on mocks with a small number of tracers (10 - 100). We compared its performance with simple mass estimators, finding that the GravSphere2 recovery is not only good, but generally superior. In particular, mass estimators can become significantly biased, while GravSphere2 not only remains unbiased, but provides constraints on the mass profile at all radii, rather than just one point at $\sim R_{1/2}$ with equal or lower errors. This suggests that, even with just 10 tracer velocities, it is still preferable to perform full GravSphere2 modeling (or at least include more flexible, non-Gaussian PDFs like the ones we use) over simpler, more commonly used approaches.
@@
While previous studies have suggested that the effectiveness of higher-order Jeans analyses is compromised without assumptions about the fourth-order anisotropy counterpart (e.g. 2013MNRAS.432.3361R; 2014MNRAS.441.1584R), we find that, not only does a general fourth-order analysis yield comparable, if not improved, constraining power to previous methods for the same data, but, in many cases, a substantially less biased recovery of the anisotropy, mass density, and slope profiles. These are all recovered at 95% confidence across almost all mocks over a wide range ($0.1\lesssim r/R_{1/2}\lesssim 10$), regardless of tracer numbers, contrasting with previous methods, including GravSphere1.5 and PyGravSphere, that are prone to notable biases, particularly at small ($R\lesssim 0.25R_{1/2}$) and large ($R\gtrsim 2R_{1/2}$) radii.
@@
With 1,000 tracers without PMs, we recover the logarithmic density slope at $R_{1/2}$ with $12\%$ (25 %) statistical errors for cuspy (cored) mock data, and with comparable absolute errors at lower radii. Including PMs, these can be distinguished with $8\%$ (12 %) errors. With only 100 tracers and no PMs, we are still able to recover slopes with $\sim 30\%$ (20%) errors. This indicates that we are able to meaningfully distinguish between cusped and cored models even with limited data, something particularly timely in light of the present interest for such searches in the literature.
@@
In summary, while PMs lead to substantially improved constraints on our mock data, LOS velocities alone are able to break the mass-anisotropy degeneracy and significantly reduce biases using our framework, establishing higher-order Jeans analysis as a viable alternative to other methods such as VSPs, which show significantly greater bias across our mock data, irrespective of tracer numbers. Improvements of our method over VSPs likely stem from some combination of biases in the VSPs themselves (see Sec. 2.1) and a full star-by-star, fourth-order treatment leading to stronger constraints that disfavor biased models (cf. Fig. 3).
@@
Finally, we contrasted GravSphere2 with classic second order Jeans modeling assuming Gaussianity. We demonstrated that, due to the mass-anisotropy degeneracy, classic Jeans models are unable to distinguish cusps from cores even with 10,000 tracers. Indeed – as has been reported previously in the literature (e.g. 2021MNRAS.501..978R) – classic Jeans can become seemingly highly biased in its recovery of the density profile, appearing to strongly favor a cusp when the mock is actually cored. We show that this bias owes to the presence of many more cuspy models than cored models in the solution hypervolume, rather than being due to cored models giving a poorer fit to the data. As such, apparent biases like this can be misleading – owing to the shape of the prior volume and unphysical model degeneracies (which GravSphere2 is able to resolve) rather than the data. We suggested a new diagnostic test to determine when posterior intervals are biased in this way: plotting the spread of models of comparable likelihood, using the 68th percentile from the maximum as a reference threshold. We showed that this diagnostic test removes the apparent bias, giving a more faithful representation of the span of models that well-represent the data.
@@
Since GravSphere2 has a distribution function, it can self-consistently model uncertain membership probabilities of individual tracers, robustly including a foreground model, which we will include in an upcoming publication Bañares-Hernández et al. (in prep.). Further additions to GravSphere2 are in preparation and will be presented in upcoming papers, including a mixture model for binary stars (following 2025arXiv250914316G) and validation of GravSphere2 on rotating mocks. Additionally, we have already incorporated in the public version of GravSphere2 the ability to include self-consistent escape velocities of stars as an additional constraint to the potential (cf. 2024Natur.631..285H).
@@
Given the present challenges concerning the structure of nearby dwarf galaxies and star clusters, GravSphere2 will be an excellent new tool for studying these systems and making the most of their ever-growing data sets. This will help us to address key outstanding questions in the field, such as whether nearby dSphs have dark matter cusps or cores (e.g. 2015MNRAS.452.3650O; 2017ARA&A..55..343B; 2019MNRAS.484.1401R; 2020JCAP...06..027K; 2020MNRAS.495...58S; 2024MNRAS.535.1015D) and whether nearby GCs host central IMBHs (e.g. 2017IJMPD..2630021M; 2020ARA&A..58..257G; 2020A&ARv..28....4N; 2023MNRAS.522.5740V; 2023arXiv231112118A; 2024Natur.631..285H; 2025A&A...693A.104B). We will apply GravSphere2 to data for nearby dSphs and GCs in forthcoming publications to address these questions.
@@ Acknowledgements.
We thank the anonymous referee for prompt and helpful comments that have helped improving this work. We are grateful to Jason Sanders for helpful comments on the use of velocity distribution functions used for this work and on the manuscript. We are grateful to Anna Genina for facilitating us access to simulated data and PyGravSphere mock data results used in this work. We would like to thank Eugene Vasiliev, Jorge Martín Camalich, Giuseppina Battaglia, José María Arroyo Polonio, Raffaele Pascale, Eduardo Vitral, and Amery Gration for helpful discussions and their interest in our work. We also thank the organizers of the “Valencia Workshop on the Small-Scale Structure of the Universe and Self-Interacting Dark Matter,” where this work was presented. The authors wish to acknowledge the contribution of the IAC High-Performance Computing support team and hardware facilities to the results of this research. ABH acknowledges funding received from the European Union through the grant “UNDARK” of the Widening participation and spreading excellence program (project number 101159929). ABH acknowledges support from the MICINN through the grant “DarkMaps” PID2022-142142NB-I00. JIR would like to acknowledge support from STFC grants ST/Y002865/1 and ST/Y002857/1. This work has made use of the following software packages: dynesty (2020MNRAS.493.3132S), corner.py (corner), NumPy (Harris_2020), SciPy (2020SciPy-NMeth), Matplotlib (Hunter:2007), Jupyter Notebook (2016ppap.book...87K), gh_alternative (2020MNRAS.499.5806S).
@@ Appendix A Gaia Challenge logarithmic density slope profile recovery with an $\alpha\beta\gamma$ mass model
In this Appendix, we consider how well GravSphere2 can recover the PlumCuspOM mock assuming an $\alpha\beta\gamma$ functional form for the enclosed mass profile rather than the default coreNFWtides model. The $\alpha\beta\gamma$ profile was used to generate the mock, and so should yield less bias. The results are shown in Fig. 8. Notice that this does indeed yield less bias as compared to using the coreNFWtides model (cf. with Fig. 4 (lower right), where the recovery is very slightly biased outside its 95% CL region).
@@ Appendix B Recovery comparison with priors
In Fig. 9 we show as an example how, even for our least constraining model of the Gaia Challenge sample – just LOS velocities of 100 tracers – the range of densities allowed by our priors (which is very similar across all the galaxies considered in our mock samples) spans several orders of magnitude and therefore has little effect on the result when using the data to constrain the mass model. We showed this by running a fit without any likelihood (just sampling the prior space) and plotting the range of densities spanned by the model in this case. Fig. 10 shows the recovery of the tracer surface density for the same mock galaxy using a binned profile with 10,000 photometric tracers. It is clear that the recovery is excellent and far in excess of the allowable range by the prior. In the cases we have considered, results will be essentially unchanged if one instead fixes the photometric profile to its best-fit value without additional marginalization. This can be used to accelerate run-times in GravSphere2.
@@ Appendix C Mass estimators with non-Gaussian PDFs and multiple tracers
Here we explore the behavior of the simple mass estimators using GravSphere2’s non-Gaussian PDFs (Sec. 2.2) for both 100 tracers and a much larger number of tracer stars (10,000). The results are shown in Fig. 11, which compares the GravSphere2 recovery of the enclosed mass for the PlumCoreOM Gaia Challenge mock with both 100 and 10,000 tracers with the same for simple mass estimators. Notice that while biases improve following these PDFs, they still persist for the Wolf and Walker estimators for 100 tracers and (to a varying extent) in all of them for the 10,000 tracer case, while GravSphere2 remains unbiased.
Figure 11: Left: GravSphere2 recovery, using GravSphere2’s non-Gaussian PDFs (Sec. 2.2), of the enclosed mass profile for the PlumCoreOM mock galaxy with only LOS velocities for 100 tracers, as compared to the same for simple mass estimators (as marked). The bands show the 95% CL values. Right: Same as left but for 10,000 tracers. While non-Gaussian PDFs improve results, a degree of bias is present in all estimators, though to a lesser extent in the Errani estimator, where the bias becomes negligible (i.e. relative to statistical errors) for 100 tracers.

### v3

@@ 4.3 Low tracer numbers and comparisons with mass estimators
Some of the most interesting objects to study at any given time are either extremely intrinsically faint or distant, meaning that they are (and perhaps always will be) scarce in tracer numbers. [^23] This should not be interpreted as prognosticating pessimism for future observations, but rather that with increasing observations it appears likely that interest will shift toward ever fainter systems that lie at the limit of the capabilities at any given time. Some current examples include the discovery of the faintest known Milky Way satellite (Smith et al., 2024b) or the dwarf galaxies in Andromeda (Collins et al., 2021; Pickett et al., 2025). In these cases, one is limited to, at most, a few tens of kinematic tracers. It is therefore of considerable interest to study the behavior of GravSphere2 in this low tracer limit, comparing it with other commonly-applied methods such as the mass estimators from Walker et al. (2009); Wolf et al. (2010); Campbell et al. (2017); Errani et al. (2018). We have done this for the case of the Fornax simulated dwarf, showing our results in Fig. 7 (left). [^24] We have omitted showing the mass estimator result from Campbell et al. (2017) as it produced a very similar result that overlapped with the one from Errani et al. (2018). For this case, we computed the global velocity dispersion by fitting GravSphere2’s non-Gaussian PDFs from Sect. 2.2 (i.e., creating one bin for all the tracers), although Gaussian PDF instead did not change our results (in this case). The errors on the estimator were then propagated from the posterior distribution of the global dispersion directly, assuming no other sources of error (systematic or otherwise). In this case, all estimators show good agreement with GravSphere2, albeit with somewhat larger errors, with all recovering the true solution within the specified 95% CL regions.
@@
Therefore, while the nominal errors of mass estimators are comparable to those of GravSphere2, the simple estimators suffer from potentially significant systematic biases that are not present in GravSphere2. We conclude that even for very low numbers of tracers (on the order of 10) it is still worth using GravSphere2 rather than these estimators, which has the additional advantage of giving constraints on the mass profile at all radii, rather than just one point at $\sim R_{1/2}$. Alternatively, particularly for more robust choices such as the Errani estimator, biases can be substantially reduced when adopting non-Gaussian PDFs, but these can still creep in as the number of tracers is increased.
We note that in our main analysis, we assumed a good knowledge of the photometric tracer profile, using our default value of 10,000 tracers for most tests so far. While it is often the case that the photometric light profile is far better constrained than the kinematic one, one may ask what the effects of restricting ourselves to a low number of photometric tracers might be on the mass estimates. Interestingly, repeating our analysis using the kinematic tracers as the only source of photometric information yielded essentially identical results. This indicates that the advantage of GravSphere2 over the mass estimators does not stem from additional information provided by a good knowledge of the tracer profile and that the uncertainty in the kinematic information dominates the uncertainty in our mass estimates.
@@ 5 Discussion and conclusion
In this work, we introduced and tested GravSphere2, an extended Jeans modeling tool that includes full treatment of the higher order Jeans equations of fourth-order moments. We introduced fourth-order PMs as a new observable to break model degeneracies, using a general radially dependent fourth-order anisotropy term as part of our analysis. This approach builds on previous works, which have only considered higher moments for LO velocities and/or made restrictive assumptions about the form of the higher order counterpart of the anisotropy.
@@
In contrast to previous versions of GravSphere, our methodology does not rely on VSPs which can become biased (see the discussion at the end of Sect. 2.1). It is also completely ”bin-free,” with a generalized velocity PDF that we used to fit discrete data star-by-star. GravSphere2 includes an improved tracer density model that has the option of fitting tracer positions individually, along with an improved sampling technique. GravSphere2 does not rely on assumptions about the form of the phase-space distribution function, as it implicitly constrains it through its velocity moments via the second- and fourth-order Jeans equations. In this way, our method is able to retain the flexibility and bias-free nature of usual Jeans analyses by not needing to specify any particular form for the distribution function, while gaining constraining power from higher order velocity information, avoiding negative moments that are unphysical (at fourth order), and offering all the other general advantages of having a distribution function.
@@
We tested these improvements using equilibrium mock data from the Gaia Challenge suite, comparing our results to those from Read et al. (2021) and Collins et al. (2021), where multiple methods were tested, including GravSphere1.5, the (until now) latest version of GravSphere, Agama, a popular distribution function-based method, and DiscreteJAM, an axisymmetric Jeans-based method. We also tested our method on simulated dwarf galaxy data recently presented in Tchiorniy and Genina (2025), which were also used to test PyGravSphere (referred to in the present paper as GravSphere1).
@@
We also tested GravSphere2 on mocks with a small number of tracers (10 - 100). We compared its performance with simple mass estimators, finding that the GravSphere2 recovery is not only good, but generally superior. In particular, mass estimators can become significantly biased, while GravSphere2 not only remains unbiased, but provides constraints on the mass profile at all radii, rather than just one point at $\sim R_{1/2}$ with equal or lower errors. This suggests that even with just ten tracer velocities, it is still preferable to perform full GravSphere2 modeling (or at least include more flexible, non-Gaussian PDFs such as the ones we use here) over simpler, more commonly used approaches.
@@
While previous studies have suggested that the effectiveness of higher order Jeans analyses is compromised without assumptions about the fourth-order anisotropy counterpart (e.g., Richardson and Fairbairn 2013, 2014), we find that, not only does a general fourth-order analysis yield comparable (if not improved) constraining power to previous methods for the same data, but, in many cases, a substantially less biased recovery of the anisotropy, mass density, and slope profiles. These are all recovered at 95% confidence across almost all mocks over a wide range ($0.1\lesssim r/R_{1/2}\lesssim 10$), regardless of tracer numbers. This stands in contrast to previous methods, including GravSphere1.5 and PyGravSphere, which are prone to notable biases, particularly at small ($R\lesssim 0.25R_{1/2}$) and large ($R\gtrsim 2R_{1/2}$) radii.
@@
Using 1,000 tracers without PMs, we recovered the logarithmic density slope at $R_{1/2}$, with $12\%$ (25 %) statistical errors for cuspy (cored) mock data and with comparable absolute errors at lower radii. When including PMs, these can be distinguished with $8\%$ (12 %) errors. With only 100 tracers and no PMs, we were still able to recover slopes with $\sim 30\%$ (20%) errors. This indicates that we are able to meaningfully distinguish between cusped and cored models even with limited data, which is particularly timely in light of the present interest for such searches in the literature.
@@
In summary, while PMs lead to substantially improved constraints on our mock data, LOS velocities alone are able to break the mass-anisotropy degeneracy and significantly reduce biases using our framework. Thus, we have established higher order Jeans analysis as a viable alternative to other methods, such as VSPs, which show significantly greater bias across our mock data, irrespective of tracer numbers. Improvements to our method over VSPs likely stem from some combination of biases in the VSPs themselves (see Sect. 2.1) and a full star-by-star, fourth-order treatment leading to stronger constraints that disfavor biased models (cf. Fig. 3).
@@
Finally, we contrasted GravSphere2 with classic second-order Jeans modeling assuming Gaussianity. We demonstrated that due to the mass-anisotropy degeneracy, classic Jeans models are unable to distinguish cusps from cores even with 10,000 tracers. Indeed, as reported previously in the literature (e.g., Read et al. 2021), classic Jeans can become seemingly highly biased in its recovery of the density profile, appearing to strongly favor a cusp when the mock is actually cored. We show that this bias owes to the presence of many more cuspy models than cored models in the solution hypervolume, rather than being due to cored models giving a poorer fit to the data. As such, apparent biases on this level can be misleading, owing to the shape of the prior volume and unphysical model degeneracies (which GravSphere2 is able to resolve), rather than the data. We suggest a new diagnostic test to determine when posterior intervals are biased in this way: plotting the spread of models of comparable likelihood, using the 68th percentile from the maximum as a reference threshold. We showed that this diagnostic test removes the apparent bias, giving a more faithful representation of the span of models that well-represent the data.
@@
Since GravSphere2 has a distribution function, it can self-consistently model uncertain membership probabilities of individual tracers, robustly including a foreground model, which we will include in an upcoming publication Bañares-Hernández et al. (in prep.). Further additions to GravSphere2 are in preparation and will be presented in upcoming papers, including a mixture model for binary stars (following Gration et al. 2025) and a validation of GravSphere2 on rotating mocks. Additionally, we have already incorporated in the public version of GravSphere2 the ability to include self-consistent escape velocities of stars as an additional constraint to the potential (cf. Häberle et al. 2024).
@@
Given the present challenges concerning the structure of nearby dwarf galaxies and star clusters, GravSphere2 will be an excellent new tool for studying these systems and making the most of their ever-growing data sets. This will help us to address key outstanding questions in the field, such as whether nearby dSphs have dark matter cusps or cores (e.g., Oman et al. 2015; Bullock and Boylan-Kolchin 2017; Read et al. 2019; Kaplinghat et al. 2020; Santos-Santos et al. 2020; De Leo et al. 2024) and whether nearby GCs host central IMBHs (e.g., Mezcua 2017; Greene et al. 2020; Neumayer et al. 2020; Vitral et al. 2023; Askar et al. 2023; Häberle et al. 2024; Bañares-Hernández et al. 2025). We will apply GravSphere2 to data for nearby dSphs and GCs in forthcoming publications to address these questions.
@@ Acknowledgements.
We thank the anonymous referee for prompt and helpful comments that have helped improving this work. We are grateful to Jason Sanders for helpful comments on the use of velocity distribution functions used for this work and on the manuscript. We are grateful to Anna Genina for facilitating us access to simulated data and PyGravSphere mock data results used in this work. We would like to thank Eugene Vasiliev, Jorge Martín Camalich, Giuseppina Battaglia, José María Arroyo Polonio, Raffaele Pascale, Eduardo Vitral, Amery Gration, and Dashuang Ye for helpful discussions and their interest in our work. We also thank the organizers of the “Valencia Workshop on the Small-Scale Structure of the Universe and Self-Interacting Dark Matter,” where this work was presented. The authors wish to acknowledge the contribution of the IAC High-Performance Computing support team and hardware facilities to the results of this research. ABH acknowledges funding received from the European Union through the grant “UNDARK” of the Widening participation and spreading excellence program (project number 101159929). ABH acknowledges support from the MICINN through the grant “DarkMaps” PID2022-142142NB-I00. JIR would like to acknowledge support from STFC grants ST/Y002865/1 and ST/Y002857/1. This work has made use of the following software packages: dynesty (Speagle, 2020), corner.py (corner), NumPy (Harris_2020), SciPy (2020SciPy-NMeth), Matplotlib (Hunter:2007), Jupyter Notebook (2016ppap.book...87K), gh_alternative (Sanders and Evans, 2020).
@@ Appendix A Gaia Challenge logarithmic density slope profile recovery with an $\alpha\beta\gamma$ mass model
In this appendix, we consider how well GravSphere2 can recover the PlumCuspOM mock assuming an $\alpha\beta\gamma$ functional form for the enclosed mass profile rather than the default coreNFWtides model. The $\alpha\beta\gamma$ profile was used to generate the mock, and so should yield less bias. The results are shown in Fig. 8. We note that this does indeed yield less bias as compared to using the coreNFWtides model (cf. with Fig. 4 (lower right), where the recovery is very slightly biased outside its 95% CL region).
@@ Appendix B Recovery comparison with priors
In Fig. 9 we show as an example how, even for our least constraining model of the Gaia Challenge sample (just the LOS velocities of 100 tracers), the range of densities allowed by our priors (which is very similar across all the galaxies considered in our mock samples) spans several orders of magnitude and therefore has little effect on the result when using the data to constrain the mass model. We showed this by running a fit without any likelihood (just sampling the prior space) and plotting the range of densities spanned by the model in this case. Fig. 10 shows the recovery of the tracer surface density for the same mock galaxy using a binned profile with 10,000 photometric tracers. It is clear that the recovery is excellent and far in excess of the allowable range by the prior. In the cases we considered, results will be essentially unchanged if one instead fixes the photometric profile to its best-fit value without additional marginalization. This can be used to accelerate run-times in GravSphere2.
@@ Appendix C Mass estimators with non-Gaussian PDFs and multiple tracers
Here, we explore the behavior of the simple mass estimators using GravSphere2’s non-Gaussian PDFs (Sect. 2.2) for both 100 tracers and a much larger number of tracer stars (10,000). The results are shown in Fig. 11, which compares the GravSphere2 recovery of the enclosed mass for the PlumCoreOM Gaia Challenge mock with both 100 and 10,000 tracers with the same for simple mass estimators. Notice that while biases improve following these PDFs, they still persist for the Wolf and Walker estimators for 100 tracers and (to a varying extent) in all of them for the 10,000 tracer case, while GravSphere2 remains unbiased.
Figure 11: Left: GravSphere2 recovery, using GravSphere2’s non-Gaussian PDFs (Sect. 2.2), of the enclosed mass profile for the PlumCoreOM mock galaxy with only LOS velocities for 100 tracers, as compared to the same for simple mass estimators (as marked). The bands show the 95% CL values. Right: Same as left but for 10,000 tracers. While non-Gaussian PDFs improve results, a degree of bias is present in all estimators, though to a lesser extent in the Errani estimator, where the bias becomes negligible (i.e., relative to statistical errors) for 100 tracers.